A 350g sample of an unknown solid is heated continuously from 10°C to 75°C. Scientists determine that the energy change during that process is 4,860J. If the solid does not change phase during the heating process, what must its specific heat capacity be?


1
Expert's answer
2021-02-22T05:56:28-0500

Solution:

q = m × C × ΔT

q = m × C × (Tf - Ti),

where:

q = amount of heat energy gained or lost by substance

m = mass of sample

C = specific heat capacity (J oC-1 g-1)

Tf = final temperature

Ti = initial temperature

 

Thus:

4860 J = (350 g) × C × (75 - 10)oC

4860 J = C × 22750 oC g

C = 0.213 J oC-1 g-1


Answer: The specific heat capacity of an unknown solid is 0.213 J oC-1 g-1.